undeadknight26 (undeadknight26):

Which shows 632 - 372 being evaluated using the difference of squares method? 632 - 372 = (3969 + 1369)(3969 - 1369) = 13,878,800 632 - 372 = 3969 - 1369 = 2,600 632 - 372 = (63 - 37)2 = 262 = 676 632 - 372 = (63 + 37)(63 - 37) = (100)(26) = 2,600

OpenStudy (solomonzelman):

difference of squares is \(\Huge\color{blue}{ \bf a^{2}-b^2=(a-b)(a+b) }\)

undeadknight26 (undeadknight26):

so b?

OpenStudy (anonymous):

is it \[(63)^2-(37)^2 ?\]

OpenStudy (solomonzelman):

All of the choices are abstruse... it's none of them -:(

undeadknight26 (undeadknight26):

yes it is 63^2 - 37^2

OpenStudy (anonymous):

then use above formula .

OpenStudy (solomonzelman):

First of all, \(\Large\color{red}{ \bf 632 - 372 = 260 }\)

undeadknight26 (undeadknight26):

it is 63^2 - 37^2...so that means its b right?

OpenStudy (solomonzelman):

Maybe you meant to say \(\Large\color{red}{ \bf 632 ^2- 372^2 ..... }\)

OpenStudy (solomonzelman):

idk, I don't see how any of the options are good. Sorry :(

undeadknight26 (undeadknight26):

no -_- i meant to say...63^2 - 37^2

OpenStudy (solomonzelman):

Oh, your question was \(\Large\color{red}{ \bf 63 ^2- 37^2 ..... }\) ?

undeadknight26 (undeadknight26):

yesh.

OpenStudy (solomonzelman):

\(\LARGE\color{green}{ \bf 63^2 - 37^2 = (63 - 37)(63+37) }\)

undeadknight26 (undeadknight26):

Thanks numba 1!

OpenStudy (solomonzelman):

Tnx for callin and for med.. g☼☺D luck !

OpenStudy (solomonzelman):

\(\LARGE\color{green}{ \bf ▓ }\)